Ronen Consulting has just realized an accounting error that has resulted in an unfunded liability of ​$380,000 due in 29 years. In other​ words, they will need ​$380,000 in 29 years. Toni​ Flanders, the​ company's CEO, is scrambling to discount the liability to the present to assist in valuing the​ firm's stock. If the appropriate discount rate is 7 ​percent, what is the present value of the​ liability?

If the appropriate discount rate is 7 ​percent, the present value of the ​$380,000 liability due in 29 is ​$ ___(Round to the nearest​ cent.)
Answer #1

We see that the present value of the liability is given as=Future Value/(1+rate)^t=380000/1.07^29=53413.86984